OpenStudy (anonymous):

The width of a rectangle is 2 ft less than the length. the area is 8ft^2. Find the length and width.

OpenStudy (anonymous):

can someone pease help

OpenStudy (anonymous):

A=L*W W=L-2 8ft^2 = L *(L-2) 8ft = L^2 -2L 0 = L^2 -2L - 8 (L-4)(L+2) length = -2 or length = 4 the length can not negative than 4 is solution checK; 8ft= 4*2 8ft= 8 length =4 , width =2

OpenStudy (anonymous):

Thank you so much